### v1

@@ Abstract
Large Language Models (LLMs) are widely used in generative applications such as chatting, code generation, and reasoning. However, many real-world workloads—such as classification, question answering, recommendation, and text embedding—rely solely on the prefill stage of inference, where the model encodes input sequences without performing autoregressive decoding. In these prefill-only scenarios, the self-attention computation becomes the primary performance bottleneck due to its quadratic complexity with respect to sequence length. In this paper, we observe that semantically different sentences often produce similar attention maps across layers and heads. Building on this insight, we propose AttnCache, a framework that accelerates the prefill stage of LLM inference by retrieving and reusing similar attention maps. Based on an attention map memoization database, AttnCache employs efficient caching and similarity search techniques to identify and reuse pre-cached attention maps during inference, thereby reducing the computational overhead of self-attention. Experimental results show that AttnCache achieves an average of 1.2 end-to-end and 2 attention speedup on CPU, and 1.6 end-to-end and 3 attention speedup on GPU, with negligible accuracy degradation. AttnCache is available at: https://github.com/dinghongsong/AttnCache.
@@ 1 Introduction
Large language models (LLMs) are extensively used in generative tasks, including chatting (e.g., ChatGPT OpenAI (2023), Deepseek DeepSeek (2025), Claude Anthropic (2025)), code generation (e.g., GitHub Copilot GitHub (2025), TraeByteDance (2025), Cursor Cursor (2025)). Each input prompt first undergoes a prefill phase to encode the context and generate the initial output token, followed by a decoding stage that autoregressively generates subsequent tokens step by step.
Nevertheless, there are also many prefill-only applications of LLMs Du et al. (2025), such as classification Wang et al. (2018); Gholamian et al. (2024); Vajjala & Shimangaud (2025), question answering Talmor et al. (2019); Hendrycks et al. (2020); Phan et al. (2025), recommendation Wang et al. (2023b); Wu et al. (2024a); Firooz et al. (2025), and data labeling He et al. (2023); Lan et al. (2024); Zhang et al. (2023a). These workloads use only the embeddings from the final layer, either to produce a single token or as input for downstream tasks. This process does not involve the decoding stage or require generating multiple tokens, so the KV cache is not reused for extended decoding. Therefore, storing the KV cache is unnecessary, and only the prefill stage of LLM inference needs to be executed. For example, in a question answering application, the input prompt could be “What is the capital of France? A. Berlin, B. London, C. Paris, D. Rome. Your answer is:”, and the LLM only needs to generate a single answer token (i.e., A, B, C, or D).
Furthermore, LLMs can also serve as text encoders to extract general-purpose sentence embeddings, excelling in text representation tasks Lee et al. (2024b); BehnamGhader et al. (2024); Lee et al. (2024a); Li et al. (2024). In such tasks, only the prefill stage is used to encode the input sentences, with either the hidden states of the last token Wang et al. (2023a); Lei et al. (2024) or the pooling of all token representations Li & Zhou (2024); Lei et al. (2025) as the sentence embedding, without involving decoding stage of LLM inference.
@@ 2 Related Work
Sentence embeddings encodes the semantic information of sentences into high-dimensional vector representations. Prior works Li & Zhou (2024); Muennighoff et al. (2024); Ni et al. (2021) have demonstrated the capability of LLMs to generate high-quality sentence embeddings. Recent studies Zhuang et al. (2024); Qin et al. (2023); Zhang et al. (2024a) have explored converting LLMs into sentence encoders without additional training. To enhance embedding quality, prompt-based techniques have gained traction. MetaEOL Lei et al. (2024) uses multitask prompts to generate general-purpose embeddings. The research by Jiang et al. (2023) illustrates how to extract a sentence embedding by prompting LLMs with the instruction “This sentence: ‘[text]’ means in one word:”. In this work, we leverage LLMs to generate sentence embeddings without fine-tuning.
@@
Most KV cache optimization approaches Beltagy et al. (2020); Zhang et al. (2023b); Oren et al. (2024) focus on accelerating the LLM decoding phase by reducing redundancy in Key and Value matrices. StreamingLLM Xiao et al. (2023) identifies “attention sinks” and keeps initial and recent tokens’ KV to anchor attention computation, while FastGen Ge et al. (2023) prunes tokens during decoding by profiling attention heads. However, these approaches do not reduce the prefill costs. In contrast, several recent efforts Wu et al. (2024b); Jiang et al. (2024); Tang et al. (2024) center on optimizing the LLM prefill phase, benefiting tasks like sentence embedding generation. PromptCache Gim et al. (2024) and ChunkAttention Ye et al. (2024) reduce time-to-first-token latency by sharing KV tensors of common prompt prefixes. Other acceleration approaches He et al. (2024); Men et al. (2024); Song et al. (2024); Zhang et al. (2024b) focus on removing the redundant attention or transformer layer in LLMs. Token pruning Ham et al. (2020); Wang et al. (2021) reduces computation by excluding less important tokens from the input, while layer-wise reuse Ying et al. (2021); Xiao et al. (2019); Bhojanapalli et al. (2021) reduces computation by sharing attention maps calculated in prior layers in multiple subsequent layers. These approaches complement AttnCache and can further enhance memory efficiency.
@@
The reuse mechanism exploits the inherent redundancy in neural networks to enhance efficiency. Prior works Ning et al. (2019); Ning & Shen (2019); Wu et al. (2022); Köpüklü et al. (2019) have explored reusing similar computation results to improve performance. Silfa et al. (2019) accelerate RNN training by reusing neuron outputs. Studies Bhojanapalli et al. (2021); Xiao et al. (2019) have shown that transformer attention maps Vaswani et al. (2017) exhibit similar distributions across adjacent layers. Many prior efforts Hunter et al. (2023); Xiao et al. (2019); Bhojanapalli et al. (2021); Ying et al. (2021); Liao & Vargas (2024) focus on sharing computed attention weights across multiple layers for the same input sequence. However, this approach may introduce dissimilar attention maps, which can degrade performance. In contrast, our work efficiently reuses similar attention maps across different sequences, overcoming the limitations of intra-sequence reuse.
@@ 3.1 Search Engine
We use two layers of Multi-Layer Perceptron (MLP) as the feature projector, which maps the input embedding to a feature vector with lower dimension size. The network structure of Feature Projector is important to the accuracy and efficiency of the search process. Compared with other embedding models, such as convolutional neural network or transformer, MLP is lightweight with less computational complexity and shorter inference time. Training the feature projector is challenging due to a lack of labeled data. Deciding the similarity between input embeddings and labeling them as similar or not is prohibitively expensive. We use the Siamese network Koch et al. (2015), which contains two identical feature projectors and shares the same weights, as shown in Figure 3.
@@
We use Smooth L1 Loss Girshick (2015) as the loss function, which is able to balance the effects of outliers. The training process iteratively updates the parameters of the feature projector to minimize the loss function.
@@ 5.1 Datasets
We take three representative datasets, including Semantic Textual Similarity (STS) Muennighoff et al. (2022) , Stanford Sentiment Treebank v2 (SST-2) Wang et al. (2018), and Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2020). STS datasets contain STS 12-16, STS-B and SICK-R Lei et al. (2024). The semantic similarity of each sentence pair is annotated with a score of 0-5. We use the Spearman correlation score Lei et al. (2024) between the ground-truth similarity scores and the predicted similarity scores as the evaluation metric. SST-2 is a binary sentiment classification dataset derived from movie reviews. Each sentence is labeled as either positive or negative. MMLU is a multi-choice dataset containing questions from 57 diverse subjects, including math, computer science, engineering, physics, and more. For SST-2 and MMLU, we use accuracy as the evaluation metric. These datasets and their outputs and labels are summarized in the Table 1.
@@ 5.2 Models
We conduct experiments using three representative open-source models, including Llama-2-7B Touvron et al. (2023), Llama-3-8B Touvron et al. (2024), and Mistral-7B Lample et al. (2023). All these models are run with weights stored in full-precision (fp32) floating-point format, and evaluation is conducted using the SentEval toolkit Conneau & Kiela (2018), measuring performance on CPU. For single-GPU scenarios, we evaluate the full-precision Llama-3-3B Touvron et al. (2024) and the 4-bit quantized Llama-3-8B, Deepseek-MoE-16B DeepSeek (2024) , and Qwen1.5-MoE-A2.7B Team (2024) on MMLU datasets. To assess performance under varying context lengths, we evaluate Llama-3-3B on MMLU under different input n-shot settings. To demonstrate the generality of AttnCache, in addition to Transformer decoder-based LLMs, we also evaluate its performance on Transformer encoder-based models, such as BERT base Devlin et al. (2018), RoBERTa Liu et al. (2019), and DeBERTa He et al. (2020).
@@ 5.3 Experimental Setting
We evaluate AttnCache on a server equipped with two sockets, each with 24-core Intel(R) Xeon(R) Silver 4410Y processors. The platform provides 512 GB DRAM and a 14 TB hard disk drive (HDD), with the DRAM used to store the attention map database and feature vector database. In addition, the platform includes an NVIDIA A100 GPU with 80 GB of HBM. Since AttnCache performs inference only once and generates sentence embeddings solely during the prefill phase—without creating or storing KV cache—a 80 GB GPU is sufficient for small models, such as Llama-3-3B. If the GPU runs out of memory, online inference can only be performed on the CPU. To build the feature vector database, we use Faiss Johnson et al. (2019), a vector database enabling efficient similarity search by the Hierarchical Navigable Small Worlds algorithm Malkov & Yashunin (2018). We use the standard LLM inference as the baseline, named full model.
@@ 5.4 Details of Implementation
For each task, we collect the input hidden states and their corresponding attention maps at each layer, which are used for training the feature projector and building databases; then we randomly select 1K samples that are not involved in training to measure AttnCache. The dimensions of the feature vector and batch size are set to 128 and 64, respectively. To maintain high inference accuracy, we set the similarity threshold to 0.99, and set , which is used to train the feature projectors (see Equation 2), to 0.2. For efficient similarity search, we construct the feature vector database using Faiss Johnson et al. (2019). Faiss is highly efficient for similarity search. For example, our evaluation shows that with Faiss, searching 100K vectors with a vector-dimension size of 128 takes less than 0.5 ms, which yields 360 and 10 speedups over self-attention computation and embedding generation, respectively. As a result, the search process does not create a performance bottleneck for AttnCache. In addition, we store each layer’s attention maps as a file object in memory. When retrieving attention maps as a batch, the file objects are mapped into a contiguous virtual memory space as a tensor without a memory copy. After self-attention calculation, the file objects are unmapped. When the combined size of the attention map and feature vector databases exceeds the available DRAM on our platform, we evaluate model performance using a hybrid storage setup that spans DRAM and HDD. While this configuration introduces I/O latency during access, it does not affect the correctness or quality of the model outputs. In this case, to evaluate the model inference time on a “virtual” big DRAM system with enough capacity to store attention maps, we use our limited DRAM assuming that the needed attention maps are in the DRAM for measuring time.
@@ 5.5 Baselines
LazyFormer Ying et al. (2021) divides all layers of the transformer to multiple subblocks. In each subblock, the attention maps are only computed in the first layer and then used by the remaining layers in the same subblock. Like LazyFormer, we set the number of layers in each sub-block to 2.
SAN Xiao et al. (2019) shares attention maps across multiple adjacent layers. But different from Lazyformer, SAN does not use a uniform subblock size (i.e., the number of transformer layers in a subblock). The subblock size is dynamically determined based on the similarity of layers in terms of the JS divergence Menéndez et al. (1997).
@@
To quantify the trade-off between speed and performance degradation, we adopt the Speedup Degradation Ratio He et al. (2024) as an evaluation metric.
@@ 5.6 Main Results
To fit the LLMs into a single GPU for MoE models, we use bitsandbytes bitsandbytes (2025) quantization to reduce the GPU memory footprint of LLMs. Specifically, for Llama-3-8B, Deepseek-MoE-16B, and Qwen1.5-MoE-A2.7B, we apply NF4 (Normal Float 4) quantization Dettmers et al. (2023). The experimental results are presented in Table 5. AttnCache achieves up to a 2.43 attention speedup and a 1.48 end-to-end speedup with only minor accuracy degradation, demonstrating the robustness of the method. These results confirm that AttnCache is both effective and generalizable across Dense and MoE model architectures.
@@ 6.1 Impacts of Model Quantinization and Pruning
Model quantization represents weights and activations with lower-precision data type, and can improve efficiency in memory usage and inference speed. We integrate AttnCache with quantization and apply Quanto Optimum (2024) to all weights, and use 4-bit quantization. We also combine AttnCache with recent LLM pruning methods, AttnDrop and BlockDrop He et al. (2024), which remove redundant attentions and layers by measuring the similarity between input and output of each layer. Table 8 shows the results. The integration of model quantization and pruning with AttnCache maintains performance: the difference between AttnCache and Quanto/BlockDrop is only 1%, and the difference between AttnCache and AttnDrop is only 2%, on average.

### v3

@@ Abstract
Large Language Models (LLMs) are widely used in generative applications such as chatting, code generation, and reasoning. However, many real-world workloads—such as classification, question answering, recommendation, and text embedding—rely solely on the prefill stage of inference, where the model encodes input sequences without performing autoregressive decoding. In these prefill-only scenarios, the self-attention computation becomes the primary performance bottleneck due to its quadratic complexity with respect to sequence length. In this paper, we observe that semantically different sentences often produce similar attention maps across layers and heads. Building on this insight, we propose AttnCache, a framework that accelerates the prefill stage of LLM inference by retrieving and reusing similar attention maps. Based on an attention map memoization database, AttnCache employs efficient caching and similarity search techniques to identify and reuse pre-cached attention maps during inference, thereby reducing the computational overhead of self-attention. Experimental results show that AttnCache achieves an average of 1.2 end-to-end and 2 attention speedup on CPU, and 1.6 end-to-end and 3 attention speedup on GPU, with negligible accuracy degradation. AttnCache is available at: https://github.com/PASAUCMerced/AttnCache
@@ 1 Introduction
Large language models (LLMs) are extensively used in generative tasks, including chatting (e.g., ChatGPT chatgpt, Deepseek deepseek2025, Claude claudeai2025), code generation (e.g., GitHub Copilot github_copilot2025, Traetraeai2025, Cursor cursor2025). Each input prompt first undergoes a prefill phase to encode the context and generate the initial output token, followed by a decoding stage that autoregressively generates subsequent tokens step by step.
Nevertheless, there are also many prefill-only applications of LLMs du2025prefillonly, such as classification wang2018glue; gholamian2024llm; vajjala2025text, question answering talmor-etal-2019-commonsenseqa; hendrycks2020measuring; phan2025humanity, recommendation wang2023enhancing; wu2024survey; firooz2025360brew, and data labeling he2023annollm; lan2024depression; zhang2023llmaaa. These workloads use only the embeddings from the final layer, either to produce a single token or as input for downstream tasks. This process does not involve the decoding stage or require generating multiple tokens, so the KV cache is not reused for extended decoding. Therefore, storing the KV cache is unnecessary, and only the prefill stage of LLM inference needs to be executed. For example, in a question answering application, the input prompt could be “What is the capital of France? A. Berlin, B. London, C. Paris, D. Rome. Your answer is:”, and the LLM only needs to generate a single answer token (i.e., A, B, C, or D).
Furthermore, LLMs can also serve as text encoders to extract general-purpose sentence embeddings, excelling in text representation tasks lee2024gecko; behnamghader2024llm2vec; lee2024nv; li2024making. In such tasks, only the prefill stage is used to encode the input sentences, with either the hidden states of the last token wang2023improving; lei2024meta or the pooling of all token representations li2024your; lei2025enhancing as the sentence embedding, without involving decoding stage of LLM inference.
@@ 2 Related Work
Sentence embeddings encodes the semantic information of sentences into high-dimensional vector representations. Prior works li2024your; muennighoff2024generative; ni2021sentence have demonstrated the capability of LLMs to generate high-quality sentence embeddings. Recent studies zhuang2024promptreps; qin2023large; zhang2024simple have explored converting LLMs into sentence encoders without additional training. To enhance embedding quality, prompt-based techniques have gained traction. MetaEOL lei2024meta uses multitask prompts to generate general-purpose embeddings. The research by jiang2023scaling illustrates how to extract a sentence embedding by prompting LLMs with the instruction “This sentence: ‘[text]’ means in one word:”. In this work, we leverage LLMs to generate sentence embeddings without fine-tuning.
@@
Most KV cache optimization approaches beltagy2020longformer; zhang2023h2o; oren2024transformers focus on accelerating the LLM decoding phase by reducing redundancy in Key and Value matrices. StreamingLLM xiao2023efficient identifies “attention sinks” and keeps initial and recent tokens’ KV to anchor attention computation, while FastGen ge2023model prunes tokens during decoding by profiling attention heads. However, these approaches do not reduce the prefill costs. In contrast, several recent efforts wu2024retrieval; jiang2024minference; tang2024razorattention center on optimizing the LLM prefill phase, benefiting tasks like sentence embedding generation. PromptCache gim2024prompt and ChunkAttention ye2024chunkattention reduce time-to-first-token latency by sharing KV tensors of common prompt prefixes. Other acceleration approaches he2024matters; men2024shortgpt; song2024sleb; zhang2024finercut focus on removing the redundant attention or transformer layer in LLMs. Token pruning ham20203; wang2021spatten reduces computation by excluding less important tokens from the input, while layer-wise reuse ying2021lazyformer; xiao2019sharing; bhojanapalli2021leveraging reduces computation by sharing attention maps calculated in prior layers in multiple subsequent layers. These approaches complement AttnCache and can further enhance memory efficiency.
@@
The reuse mechanism exploits the inherent redundancy in neural networks to enhance efficiency. Prior works ning2019adaptive; ning2019deep; wu2022drew; kopuklu2019convolutional have explored reusing similar computation results to improve performance. silfa2019neuron accelerate RNN training by reusing neuron outputs. Studies bhojanapalli2021leveraging; xiao2019sharing have shown that transformer attention maps vaswani2017attention exhibit similar distributions across adjacent layers. Many prior efforts hunter2023fast; xiao2019sharing; bhojanapalli2021leveraging; ying2021lazyformer; liao2024beyond focus on sharing computed attention weights across multiple layers for the same input sequence. However, this approach may introduce dissimilar attention maps, which can degrade performance. In contrast, our work efficiently reuses similar attention maps across different sequences, overcoming the limitations of intra-sequence reuse.
@@ 3.1 Search Engine
We use two layers of Multi-Layer Perceptron (MLP) as the feature projector, which maps the input embedding to a feature vector with lower dimension size. The network structure of Feature Projector is important to the accuracy and efficiency of the search process. Compared with other embedding models, such as convolutional neural network or transformer, MLP is lightweight with less computational complexity and shorter inference time. Training the feature projector is challenging due to a lack of labeled data. Deciding the similarity between input embeddings and labeling them as similar or not is prohibitively expensive. We use the Siamese network koch2015siamese, which contains two identical feature projectors and shares the same weights, as shown in Figure 3.
@@
We use Smooth L1 Loss girshick2015fast as the loss function, which is able to balance the effects of outliers. The training process iteratively updates the parameters of the feature projector to minimize the loss function.
@@ 5.1 Datasets
We take three representative datasets, including Semantic Textual Similarity (STS) muennighoff2022mteb , Stanford Sentiment Treebank v2 (SST-2) wang2018glue, and Massive Multitask Language Understanding (MMLU) hendrycks2020measuring. STS datasets contain STS 12-16, STS-B and SICK-R lei2024meta. The semantic similarity of each sentence pair is annotated with a score of 0-5. We use the Spearman correlation score lei2024meta between the ground-truth similarity scores and the predicted similarity scores as the evaluation metric. SST-2 is a binary sentiment classification dataset derived from movie reviews. Each sentence is labeled as either positive or negative. MMLU is a multi-choice dataset containing questions from 57 diverse subjects, including math, computer science, engineering, physics, and more. For SST-2 and MMLU, we use accuracy as the evaluation metric. These datasets and their outputs and labels are summarized in the Table 1.
@@ 5.2 Models
We conduct experiments using three representative open-source models, including Llama-2-7B touvron2023llama, Llama-3-8B touvron2024llama3, and Mistral-7B jiang2023mistral. All these models are run with weights stored in full-precision (fp32) floating-point format, and evaluation is conducted using the SentEval toolkit conneau2018senteval, measuring performance on CPU. For single-GPU scenarios, we evaluate the full-precision Llama-3-3B touvron2024llama3 and the 4-bit quantized Llama-3-8B, Deepseek-MoE-16B deepseekmoe2024 , and Qwen1.5-MoE-A2.7B qwen2024qwen on MMLU datasets. To assess performance under varying context lengths, we evaluate Llama-3-3B on MMLU under different input n-shot settings. To demonstrate the generality of AttnCache, in addition to Transformer decoder-based LLMs, we also evaluate its performance on Transformer encoder-based models, such as BERT base devlin2018bert, RoBERTa liu2019roberta, and DeBERTa he2020deberta.
@@ 5.3 Experimental Setting
We evaluate AttnCache on a server equipped with two sockets, each with 24-core Intel(R) Xeon(R) Silver 4410Y processors. The platform provides 512 GB DRAM and a 14 TB hard disk drive (HDD), with the DRAM used to store the attention map database and feature vector database. In addition, the platform includes an NVIDIA A100 GPU with 80 GB of HBM. Since AttnCache performs inference only once and generates sentence embeddings solely during the prefill phase—without creating or storing KV cache—a 80 GB GPU is sufficient for small models, such as Llama-3-3B. If the GPU runs out of memory, online inference can only be performed on the CPU. To build the feature vector database, we use Faiss johnson2019billion, a vector database enabling efficient similarity search by the Hierarchical Navigable Small Worlds algorithm malkov2018efficient. We use the standard LLM inference as the baseline, named full model.
@@ 5.4 Details of Implementation
For each task, we collect the input hidden states and their corresponding attention maps at each layer, which are used for training the feature projector and building databases; then we randomly select 1K samples that are not involved in training to measure AttnCache. The dimensions of the feature vector and batch size are set to 128 and 64, respectively. To maintain high inference accuracy, we set the similarity threshold to 0.99, and set , which is used to train the feature projectors (see Equation 2), to 0.2. For efficient similarity search, we construct the feature vector database using Faiss johnson2019billion. Faiss is highly efficient for similarity search. For example, our evaluation shows that with Faiss, searching 100K vectors with a vector-dimension size of 128 takes less than 0.5 ms, which yields 360 and 10 speedups over self-attention computation and embedding generation, respectively. As a result, the search process does not create a performance bottleneck for AttnCache. In addition, we store each layer’s attention maps as a file object in memory. When retrieving attention maps as a batch, the file objects are mapped into a contiguous virtual memory space as a tensor without a memory copy. After self-attention calculation, the file objects are unmapped. When the combined size of the attention map and feature vector databases exceeds the available DRAM on our platform, we evaluate model performance using a hybrid storage setup that spans DRAM and HDD. While this configuration introduces I/O latency during access, it does not affect the correctness or quality of the model outputs. In this case, to evaluate the model inference time on a “virtual” big DRAM system with enough capacity to store attention maps, we use our limited DRAM assuming that the needed attention maps are in the DRAM for measuring time.
@@ 5.5 Baselines
LazyFormer ying2021lazyformer divides all layers of the transformer to multiple subblocks. In each subblock, the attention maps are only computed in the first layer and then used by the remaining layers in the same subblock. Like LazyFormer, we set the number of layers in each sub-block to 2.
SAN xiao2019sharing shares attention maps across multiple adjacent layers. But different from Lazyformer, SAN does not use a uniform subblock size (i.e., the number of transformer layers in a subblock). The subblock size is dynamically determined based on the similarity of layers in terms of the JS divergence menendez1997jensen.
@@
To quantify the trade-off between speed and performance degradation, we adopt the Speedup Degradation Ratio he2024matters as an evaluation metric.
@@ 5.6 Main Results
To fit the LLMs into a single GPU for MoE models, we use bitsandbytes bitsandbytes2025 quantization to reduce the GPU memory footprint of LLMs. Specifically, for Llama-3-8B, Deepseek-MoE-16B, and Qwen1.5-MoE-A2.7B, we apply NF4 (Normal Float 4) quantization dettmers2023qlora. The experimental results are presented in Table 5. AttnCache achieves up to a 2.43 attention speedup and a 1.48 end-to-end speedup with only minor accuracy degradation, demonstrating the robustness of the method. These results confirm that AttnCache is both effective and generalizable across Dense and MoE model architectures.
@@ 6.1 Impacts of Model Quantinization and Pruning
Model quantization represents weights and activations with lower-precision data type, and can improve efficiency in memory usage and inference speed. We integrate AttnCache with quantization and apply Quanto Optimum-quanto to all weights, and use 4-bit quantization. We also combine AttnCache with recent LLM pruning methods, AttnDrop and BlockDrop he2024matters, which remove redundant attentions and layers by measuring the similarity between input and output of each layer. Table 8 shows the results. The integration of model quantization and pruning with AttnCache maintains performance: the difference between AttnCache and Quanto/BlockDrop is only 1%, and the difference between AttnCache and AttnDrop is only 2%, on average.